# Enhancing Weather Predictions: Super-Resolution via Deep Diffusion Models

### v1

@@ 3.1 SR3
$q\left(y_{t}\mid y_{0}\right)=\mathcal{N}\left(y_{t}|\sqrt{\bar{\alpha}_{t}}y_{0},(1-\bar{\alpha}_{t})I\right)$
@@
In the reverse process, the goal is to reconstruct the original image $x_{0}$ from the noise, where the distribution of noise is $\mathcal{N}(\mathbf{0},\boldsymbol{I})$. Direct computation of $q(x_{t}|x_{t-1})$ is impossible; instead, reconstruction is approximated by a learned model $p_{\theta}(x_{t-1}|x_{t})$ with parameters $\theta$ where $x$ is source image:
${p}_{{\theta}}{\left({y}_{{{t}-{1}}}\mid{y}_{{{t}}},{x}\right)}={\mathcal{N}}{\left({y}_{{{t}-{1}}}\mid\mu_{{\theta}}{\left({x},{y}_{{{t}}},\bar{\alpha}_{t}\right)},{\sigma_{{{t}}}^{{{2}}}}{I}\right)}$
@@ 3.2 ResDiff
Furthermore, ResDiff introduces High-Frequency Guided Diffusion that modifies the diffusion model to focus on and enhance high-frequency image components. This method ensures that the details lost in the SR3 diffusion process, are meticulously reconstructed. As shown in Figure 1, HF Guided Diffusion adds the FD Info Splitter and HF-guided Cross-Attention (CA) into the architecture.
@@ 3.2.1 FD Info Splitter
The Frequency-Domain Information Splitter (FD Info Splitter) is a novel component of ResDiff that segregates image data into high and low-frequency bands. This segregation is fundamental for focusing the model’s attention on preserving or enhancing the details critical for high-quality image super-resolution. The process begins with the application of a 2D Fast Fourier Transform (FFT) to both the interpolated and the noised images. This transformation allows the model to analyze and manipulate the frequency components of the image data directly.
@@
where $M$ represents the feature map obtained from the FFT, and $l$ denotes a predetermined upper limit for the filter’s standard deviation. This calculation ensures that the emphasis on high-frequency details is adaptively scaled, enhancing the model’s focus on areas of the image requiring finer detail reconstruction.
The high-pass filtering, represented by $H(u,v)$, is then applied to isolate and enhance these high-frequency components:
@@
where $D(u,v)$ measures the distance of a frequency component from the origin in the frequency domain. This filtering process results in a feature map $L$, enriched with high-frequency details.
@@
The culmination of this process produces a set of five distinct feature maps: $[x,y,x_{HF},x_{LF},x^{\prime}_{t}]$, where $x$ and $y$ represent the source and the noise-added images, respectively, $x_{HF}$ and $x_{LF}$ denote the high and low-frequency components extracted, and $x^{\prime}_{t}$ is the target high-resolution output at iteration $t$. These feature maps are integral to the ResDiff model’s ability to reconstruct images with enhanced detail and clarity, particularly in the high-frequency bands that are crucial for perceptual quality.
3.2.2 HF-guided Cross-Attention (CA)
The HF-guided Cross-Attention mechanism further refines the model’s capacity to focus on and enhance high-frequency details. By leveraging the high-frequency component obtained by Discrete wavelet transformation [7], this mechanism directs the model’s attention specifically towards areas of the image that benefit most from detail enhancement. This targeted approach ensures that the diffusion process is finely tuned to the nuances of image textures and edges, significantly improving the reconstruction of details lost in the initial SR3 diffusion stages.
@@ 5 Results
Figure 3: Model validation scores during training
@@
Figure 4: In the left column, the HR reference image and images generated by the models are displayed, annotated with their corresponding temperatures in Kelvin. The right column shows the difference (delta) between HR reference and generated super-resolution images.

### v2

@@ 3.1 SR3
$q\left(y_{t}\mid y_{0}\right)=\mathcal{N}\left(y_{t};\sqrt{\bar{\alpha}_{t}}y_{0},(1-\bar{\alpha}_{t})I\right)$
@@
In the reverse process, the goal is to reconstruct the original image $y_{0}$ from the noise, where the distribution of noise is $\mathcal{N}(\mathbf{0},\boldsymbol{I})$. Direct computation of $q(x_{t}|x_{t-1})$ is impossible; instead, reconstruction is approximated by a learned model $p_{\theta}(x_{t-1}|x_{t})$ with parameters $\theta$ where $x$ is low resolution source image:
${p}_{{\theta}}{\left({y}_{{{t}-{1}}}\mid{y}_{{{t}}},{x}\right)}={\mathcal{N}}{\left({y}_{{{t}-{1}}};\mu_{{\theta}}{\left({x},{y}_{{{t}}},\bar{\alpha}_{t}\right)},{\sigma_{{{t}}}^{{{2}}}}{I}\right)}$
@@ 3.2 ResDiff
Furthermore, ResDiff introduces High-Frequency Guided Diffusion that modifies the diffusion model to focus on and enhance high-frequency image components. This method ensures that the details lost in the SR3 diffusion process, are meticulously reconstructed. As shown in Figure 1, HF Guided Diffusion adds the FD Info Splitter and HF-guided Cross-Attention into the architecture.
@@ 3.2.1 FD Info Splitter
The Frequency-Domain Information Splitter (FD Info Splitter) [11] is a novel component of ResDiff that segregates image data into high and low-frequency bands. This segregation is fundamental for focusing the model’s attention on preserving or enhancing the details critical for high-quality image super-resolution. The process begins with the application of a 2D Fast Fourier Transform (FFT) to both the interpolated and the noised images. This transformation allows the model to analyze and manipulate the frequency components of the image data directly.
@@
where $M$ represents the feature map obtained from the FFT, and $l$ denotes a predetermined upper limit for the filter’s standard deviation. This calculation ensures that the emphasis on high-frequency details is adaptively scaled, enhancing the model’s focus on areas of the image requiring finer detail reconstruction. The high-pass filtering, represented by $H(u,v)$, is then applied to isolate and enhance these high-frequency components:
@@
where $D(u,v)$ measures the distance of a frequency component from the origin in the frequency domain. This filtering process results in a feature map $L$, enriched with high-frequency details. [11]
@@
The culmination of this process produces a set of five distinct feature maps: $[x,y,x_{HF},x_{LF},x^{\prime}_{t}]$, where $x$ and $y$ represent the source and the noise-added images, respectively, $x_{HF}$ and $x_{LF}$ denote the high and low-frequency components extracted, and $x^{\prime}_{t}$ is the target high-resolution output at iteration $t$. These feature maps are integral to the ResDiff model’s ability to reconstruct images with enhanced detail and clarity, particularly in the high-frequency bands that are crucial for perceptual quality [11].
3.2.2 HF-guided Cross-Attention
The HF-guided Cross-Attention [11] mechanism further refines the model’s capacity to focus on and enhance high-frequency details. By leveraging the high-frequency component obtained by Discrete wavelet transformation [7], this mechanism directs the model’s attention specifically towards areas of the image that benefit most from detail enhancement. This targeted approach ensures that the diffusion process is finely tuned to the nuances of image textures and edges, significantly improving the reconstruction of details lost in the initial SR3 diffusion stages.
@@ 5 Results
Figure 3: Validation scores during training across models: SR3, ResDiff, ResDiff + Physics.
@@
Figure 4: In the left column, the HR reference image and images generated by the models are displayed, annotated with their corresponding temperatures in Kelvin. The right column shows the absolute error between HR reference and super-resolution images.